Determine whether or not the following vectors span
u1 = ( 1, 1, 2), u2 = ( 1, -1, 2 ) and u3 = ( 1, 0, 1 ).
1
Expert's answer
2020-12-16T20:02:12-0500

Taking into account that "\left|\begin{array}{ccc} 1 & 1 & 2 \\ 1 & -1 & 2 \\ 1 & 0 & 1\end{array}\right |= -1+2-(-2+1)=1+1=2\ne 0," we conclude that the vectors "u_1 = ( 1, 1, 2), u_2 = ( 1, -1, 2 )" and "u_3 = ( 1, 0, 1 )" are linearly independent, and consequently span "\mathbb R^3."